The temperature increases from 6:00 A.M. onward. Let the input t represent the time of the day, and let the output F represent the temperature in degrees Fahrenheit. At 6:00 A.M. (t = 6), the temperature is F = 55. At 10:00 A.M. (t = 10), the temperature is F = 69. Find the linear relationship between t and F.
Answer by Boreal(15235)

those points are (6, 55) and (10, 69)
the slope is change in y/change in x = 14/4=3.5
point slope formula y-y1=m(x-x1) m slope, (x1, y1) point
y-55=3.5(x-6)
y=3.5x-21+55
y=3.5x +34